OpenStudy (anonymous):

Type the second number (in decimal form to tenths) for each pair such that square root is 6 is the geometric mean for the pair. What is the second number in the pair if the first number is 12?

OpenStudy (anonymous):

.6

OpenStudy (watchmath):

We want to find x such that \(\sqrt{12x}=\sqrt{6}\) Hence \(x=1/2\).

OpenStudy (watchmath):

In decimal form, x=0.5